Your textbook says that the focal length f of a lens is given by the lensmaker's formula 1/f=0.6(1/r1+1/r2) where r1 and r2 are the radii of the two circular surfaces. If those radii are 4 and 6 respectively, find f.
A. 3/25
B. 4
C. 16/67
D. 1/4

Found 2 solutions by marksmartuko, Alan3354:
Answer by marksmartuko(1)
Answer by Alan3354(69443)

Your textbook says that the focal length f of a lens is given by the lensmaker's formula 1/f=0.6(1/r1+1/r2) where r1 and r2 are the radii of the two circular surfaces. If those radii are 4 and 6 respectively, find f.
-------------
1/f=0.6(1/r1+1/r2)
Sub for r1 & r1
1/f=0.6(1/4 + 1/6)
Solve for f